# Isospin asymmetry and neutron stars in V-QCD

### v1

@@ 4.2 General configuration: $\beta$-equilibrated neutral matter
Figure 3: Particle fractions as functions of density in units of the saturation density. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}$. For the color coding, see fig. 2. Thick lines represent proton fractions, while thin lines (always below the thick ones) represent electron fractions: Muon fractions can be easily derived as the difference between the two, as dictated by charge neutrality, and neutron fractions are simply given by subtracting the proton fractions from unity. For every plot we also show in purple the proton fractions predicted by the SLy4 EOS at densities lower than the matching density.
@@
where $\mu_{N}$, $\mu_{P}$ are the chemical potentials of the neutron and the proton, respectively , and $n_{l}$ is the lepton number density. The leptonic free energy is that of a (massive) Fermi gas:
@@
Having now a family of solutions determined by the single parameter [^6] Note, however, that eq. (51) is derived assuming that the parameters $n_{B}$ and $n_{I}$ are read off of the asymptotics of the $A_{t}$ field rescaled with the overall coefficient of the action, as per standard holographic dictionary: This implies that for matching procedure (a), they both get rescaled with $c_{b}$. Because of this, the equation of state will now depend from $c_{b}$ in a nontrivial way: Before the introduction of $\beta$-equilibrium, it amounts to an overall rescaling of density, energy and pressure. We are considering now to have already fixed $c_{b}$, $b_{1}$ to the values appropriate for the matching, while practically the procedure involves solving eq. (51) at a fixed matching density for many values of $(b_{1},c_{b})$ and determining the pairs that produce a continuous equation of state after matching. $n_{B}$, we can proceed to compute the equation of state.
@@
$\Omega_{b}=-\left(S_{TDBI}^{on-shell}+S_{TCS}^{on-shell}\right).$ (52)
@@
While this is sufficient to compute the EOS for the baryonic phase above the transition density $n_{t}$, it does not account yet for the possibility of a quark matter phase, which will be discussed in the next section.
@@ 5 Quark phase
The onset of quark matter can be a decisive factor in establishing the highest mass a neutron star can achieve. If we want to compare our result with observations, we need to account for the possibility of quark matter influencing the MR curves of neutron stars. To model the dense quark-matter phase, we follow Alho:2012mh ; Alho:2013hsa and compute the equation of state from the thermodynamics of charged black hole solutions in V-QCD. For this computation, we need to include the effects of backreaction of the gauge fields onto the geometry: To turn on a nontopological charge density, a horizon in the background is needed to impose Dirichlet boundary conditions on the Abelian vectorial gauge field $\widehat{A}_{t}$. Other components of the gauge field are set to zero. For the metric, we take the Ansatz of eq. (5) but instead of setting the blackening factor $f(r)$ to one it is now a free function. We find the solution that describes a geometry with a planar horizon ($f(r_{h})=0$ for some value of $r=r_{h}$) by numerically solving the equations of motion arising from the gluon and flavor actions in eqs. (3) and (6). For the potentials we are using, only a “hairless” chirally symmetric black hole phase solution appears, so that the tachyon field $T$ vanishes. This solution is dual to a chirally symmetric quark-gluon plasma.
At zero quark masses, and accounting for the presence of the strange quark, $\beta$-equilibrium is trivially solved by an isospin symmetric configuration $n_{u}=n_{d}=n_{s}$, setting $\mu_{I}=0$ and similarly having vanishing lepton number densities. This is of course an approximation, in reality we should account for quark masses, including the larger strange mass. The onset of the quark phase is then uniquely determined by requiring continuity of the pressure and baryon number chemical potential $\mu_{B}$ at the transition, while the baryon number density is in general discontinuous. We follow the approach Ishii:2019gta ; Ecker:2019xrw ; Jokela:2020piw where the pressure from nuclear matter is directly compared to the pressure of quark matter (or, more precisely, the pressure difference between the quark matter and the vacuum) despite the fact that the nuclear matter pressure was computed in the probe approximation whereas the quark matter solution is fully backreacted.
@@
Figure 4: Collection of equations of state. Top left panel: Matching procedure (b), the black part of the curves is derived from V-QCD and coincides for all the choices of low density EOS. Remaining panels: Matching procedure (a) with matching density $n_{t}$. For the color coding, see fig. 2. For each solid (dashed) curve and for each color, a colored dot (diamond) corresponds to the central pressure and energy density of the heaviest neutron star that the EOS supports.
@@
Figure 5: Speed of sound squared as a function of baryonic density. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}$. The horizontal gap in these speeds of sound manifests due to the phase transition from baryonic to quark matter. The vertical lines are displayed as guides for the beginning and ending of the gaps. For the color coding, see fig. 2. The horizontal black line marks the conformal value $c_{s}^{2}=1/3$ for reference.
@@
In fig. 5, we display the speed of sound squared for all the EOSs computed, but taking into account also the transition to the quark matter phase, which manifests itself as a horizontal gap in the plots. Note that we could only take the quark matter phase transition into account for the procedure (a).
@@ 6 Neutron stars
The TOV equations can be solved by choosing initial values $P(r=0)=P_{0},m(r=0)=0$ and identifying the value $R$ at which $P(R)=0$ as the radius of the star (correspondingly, the mass of the star is given by $M=m(R)$). Repeating the process for all values of $P_{0}$ that result in stable stars leads to a curve in the $M$-$R$ plane describing all possible neutron stars for each given EOS.
@@
Figure 6: Mass-Radius (MR) curves for neutron stars. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}$. For the color coding, see fig. 2.
Our results for this procedure are shown in fig. 6, where we also report the constraints from NICER Riley:2019yda ; Miller:2019cac ; Riley:2021pdl ; Miller:2021qha and an estimate on the highest possible mass as determined in ref. Margalit:2017dij . Almost all of our EOSs are stiff enough to reach the range of two solar masses as required, for example, the NICER measurement of the mass of J0740+6620. However, the EOSs with $n_{t}=1.8n_{S}$ and the soft variant of the HLPS EOSs are so soft that they are in tension with this measurement. Moreover, many other EOSs, including all EOSs constructed with the procedure (b), lead to slightly larger maximum masses than the bound of Margalit:2017dij . Note however that this bound is a rather indirect result of the analysis of the GW170817 binary merger event.
@@
Figure 7: Mass-Tidal Deformability (MTD) curves for neutron stars. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}$. For the color coding, see fig. 2.
@@
Figure 8: Density as a function of radial distance from the center of neutron stars, for typical stars of mass $1.4M_{\odot}$. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}$. For the color coding, see fig. 2. Thick lines, thin lines, and darker thick lines correspond respectively to the holographic core, the low density core (HPLS or SLy4) and the crust of the stars. Note that in the bottom-right panel, only one red curve appears; this is because for this star the density never reaches the matching density $n_{t}=0.288\rm{fm}^{-3}$, and so it is insensitive to the holographic model.
@@ 7 Conclusions
With this work, we have studied the effects of isospin breaking in nuclear matter within the holographic V-QCD model. A crucial simplifying approximation utilized in this work is the homogeneous Ansatz for nuclear matter, which consists of introducing a finite baryonic density via a discontinuity in the bulk gauge fields, which are taken to be independent of the position in three dimensional space. Isospin is introduced via an external chemical potential, following the usual holographic prescription of encoding its value in the UV boundary condition for the field $A^{a=3}_{t}$. The approximation inherent in the homogeneous Ansatz is expected to be reliable only at high density: For this reason, we used a set of phenomenological equations of state for the description of the low-density regime (SLy4 and three representative HLPS). We considered two different approaches to matching the low density EOS with the holographic high density ones: The one labeled “procedure (a)” is more phenomenology-driven and corresponds to the introduction of a new parameter that acts as a rescaling factor for the holographic action, while with the one labeled “procedure (b)” avoid introducing new parameters and instead fit the already present ones to correctly reproduce the value of saturation density.
@@
Finally, having the computed EOS in hand, we solved the TOV equations and derived the properties of static neutron stars. We computed MR curves and tidal deformabilities of the stars resulting from each EOS, finding results that are in all cases close to observations, and a subset of EOSs that satisfy the constraints from NICER on the MR curves and from LIGO/Virgo on the tidal deformability. An interesting result is that those EOSs that produce neutron stars that satisfy these bounds are also those that correspond to parameters that correctly reproduced the phenomenological symmetry energy at saturation: This may hint that the rescaling procedure (a) tends to successfully capture some physics that is lost when taking the approximations of homogeneity and of the flat-space dilaton potential.
These and other improvements are for now left for future works.
Acknowledgments
Appendix A Vanishing contribution of variation of $r_{c}$ to the symmetry energy
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{2}\mu_{I}^{2},$ (59)
$\mathcal{L}_{0}(r_{c})+\mathcal{L}_{2}(r_{c})\mu_{I}^{2}=0.$ (60)
We can then expand also the critical radius $r_{c}$ as:
$r_{c}=r_{c0}+\delta r_{c},$ (61)
with $r_{c0}$ being the critical position of the discontinuity for symmetric matter, and $\delta r_{c}$ being its $\mu_{I}$-dependent infinitesimal variation. Imposing the equilibrium condition at the unperturbed order allows us to determine $\delta(r_{c})$, which is also quadratic in $\mu_{I}$:
Appendix B Tables of parameters
In tab. 1 we report the choices of values of the parameters $n_{t},b_{1},c_{b}$ for each hybrid equation of state. On the left are the choices for potentials 7a, while on the right for potentials 5b. In the last column labeled “Data” we evaluate the compatibility of the resulting neutron stars with the NICER and LIGO/Virgo data, indicating with a check mark ✓ full compatibility, while with ✗ the presence of tension with the data currently available.
@@ Appendix B Tables of parameters
Table 1: Summary of EOS and parameters for matching procedure (a).

### v2

@@ 4.2 General configuration: $\beta$-equilibrated neutral matter
Figure 3: Proton (thick lines) and electron fractions (thin lines) as functions of baryon number density, $n_{B}$, in units of the saturation density $n_{S}$. The muon fraction is inferred from charge neutrality as the difference between the electron and proton fractions, and the neutron fraction is obtained by subtracting the proton fraction from unity. Color coding denotes the low-density equations of state used for matching: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4. Solid (dashed) lines correspond to the V-QCD potential 7a (5b). The top-left panel displays results from matching procedure (b), whereas the remaining panels correspond to procedure (a) for transition densities $n_{t}=\{1.2,1.5,1.8\}n_{S}$. Purple curves show the SLy4 predictions at densities below the matching point for reference.
@@
where $\mu_{N}$ and $\mu_{P}$ are the chemical potentials of the neutron and the proton, respectively , and $n_{l}$ is the lepton number density. The leptonic free energy is that of a (massive) Fermi gas:
@@
Having now a family of solutions determined by the single parameter [^6] Note, however, that eq. (51) is derived assuming that the parameters $n_{B}$ and $n_{I}$ are read off of the asymptotics of the $A_{t}$ field rescaled with the overall coefficient of the action, as per standard holographic dictionary: This implies that for matching procedure (a), they both get rescaled with $c_{b}$. Because of this, the equation of state will now depend on $c_{b}$ in a nontrivial way: Before the introduction of $\beta$-equilibrium, it amounts to an overall rescaling of density, energy and pressure. We are considering now to have already fixed $c_{b}$, $b_{1}$ to the values appropriate for the matching, while practically the procedure involves solving eq. (51) at a fixed matching density for many values of $(b_{1},c_{b})$ and determining the pairs that produce a continuous equation of state after matching. $n_{B}$, we can proceed to compute the equation of state.
@@
$\Omega_{b}=-\left(S_{\rm TDBI}^{\rm on-shell}+S_{\rm TCS}^{\rm on-shell}\right).$ (52)
@@
While the calculation up to now is sufficient for computing the EOS for the baryonic phase above the transition (matching) density $n_{t}$, it does not account yet for the possibility of a quark matter phase, which will be discussed in the next section.
@@ 5 Quark phase
The onset of quark matter can be a decisive factor in establishing the highest mass a neutron star can achieve. If we want to compare our results with observations, we need to account for the possibility of quark matter influencing the MR curves of neutron stars. To model the dense quark-matter phase, we follow Alho:2012mh ; Alho:2013hsa and compute the equation of state from the thermodynamics of charged black hole solutions in V-QCD. For this computation, we need to include the effects of backreaction of the gauge fields onto the geometry: To turn on a nontopological charge density, a horizon in the background is needed to impose Dirichlet boundary conditions on the Abelian vectorial gauge field $\widehat{A}_{t}$. Other components of the gauge field are set to zero. For the metric, we take the Ansatz of eq. (5) but instead of setting the blackening factor $f(r)$ to one it is now a free function. We find the solution that describes a geometry with a planar horizon ($f(r_{h})=0$ for some value of $r=r_{h}$) by numerically solving the equations of motion arising from the gluon and flavor actions in eqs. (3) and (6). For the potentials we are using, only a “hairless” chirally symmetric black hole phase solution appears, so that the tachyon field $T$ vanishes. This solution is dual to a chirally symmetric quark-gluon plasma.
At zero quark masses, and accounting for the presence of the strange quark, $\beta$-equilibrium is trivially solved by an isospin symmetric configuration $n_{u}=n_{d}=n_{s}$, setting $\mu_{I}=0$ and similarly having vanishing lepton number densities. This is of course an approximation, in reality we should account for quark masses, including the larger strange mass. The onset of the quark phase is then uniquely determined by requiring continuity of the pressure and the baryon number chemical potential $\mu_{B}$ at the transition, while the baryon number density is in general discontinuous. We follow the approach Ishii:2019gta ; Ecker:2019xrw ; Jokela:2020piw where the pressure from nuclear matter is directly compared to the pressure of quark matter (or, more precisely, the pressure difference between the quark matter and the vacuum) despite the fact that the nuclear matter pressure was computed in the probe approximation whereas the quark matter solution is fully backreacted.
@@
Figure 4: Collection of equations of state. Color coding and line styles are as in fig. 2: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4; solid (dashed) lines correspond to potential 7a (5b). The top-left panel corresponds to procedure (b), with the black segments representing the V-QCD high-density part. The other panels correspond to procedure (a) with transition densities $n_{t}=\{1.2,1.5,1.8\}n_{S}$. Colored dots (diamonds) mark the central energy densities and pressures of the most massive neutron stars supported by each EOS. The gray band is obtained from all the quadrutropic EOS interpolating from EFT results (low energy density) to perturbative QCD ones (high energy density) by requiring the TOV mass $M_{\rm TOV}$ and the tidal deformability $\Lambda$ to satisfy $M_{\rm TOV}>2M_{\odot}$ and $\Lambda_{1.4M_{\odot}}<580$ Jarvinen:2021jbd .
@@
Figure 5: Speed of sound squared as a function of the baryon number density for each hybrid EOS. The black horizontal line corresponds to the conformal limit $c_{s}^{2}=1/3$. Color coding and line styles are as in fig. 2: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4; solid (dashed) lines correspond to potential 7a (5b). The top-left panel refers to matching procedure (b), while the others correspond to procedure (a) with transition densities $n_{t}=\{1.2,1.5,1.8\}n_{S}$. Horizontal gaps indicate the first-order phase transition between baryonic and quark matter phases, bounded by vertical guide lines.
@@
In fig. 5, we display the speed of sound squared for all the EOSs computed, but take into account also the transition to the quark matter phase, which manifests itself as a horizontal gap in the plots. Note that we could only take the quark matter phase transition into account for the procedure (a).
@@ 6 Neutron stars
The TOV equations are solved with initial values $P(r=0)=P_{0}$, $m(r=0)=0$ and the value $R$ is determined from $P(R)=0$, which is identified as the radius of the star (correspondingly, the mass of the star is given by $M=m(R)$). Repeating the process for all values of $P_{0}$ that result in stable stars leads to a curve in the $M$-$R$ plane describing all possible neutron stars for each given EOS.
@@
Figure 6: Mass-radius (MR) relations for neutron stars obtained by solving the TOV equations with the constructed EOS. Each curve represents a different hybrid EOS, with color coding and line styles as in fig. 2: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4; solid (dashed) lines correspond to potential 7a (5b). The top-left panel corresponds to matching procedure (b), while the others show procedure (a) with $n_{t}=\{1.2,1.5,1.8\}n_{S}$. The gray shaded region marks observational constraints from NICER, and the blue shaded area indicates an estimate for the highest possible mass value $M_{\rm TOV}$ for a static neutron star, obtained combining results from NICER and the analysis of GW170817.
Our results for this procedure are shown in fig. 6, where we also report the constraints from NICER Riley:2019yda ; Miller:2019cac ; Riley:2021pdl ; Miller:2021qha and an estimate on the highest possible mass as determined in ref. Margalit:2017dij . Almost all of our EOSs are stiff enough to reach the range of two solar masses as required, for example, by the NICER measurement of the mass of J0740+6620. However, the EOSs with $n_{t}=1.8n_{S}$ and the soft variant of the HLPS EOSs are so soft that they are in tension with this measurement. Moreover, many other EOSs, including all EOSs constructed with the procedure (b), lead to slightly larger maximum masses than the bound of Margalit:2017dij . Note however that this bound is a rather indirect result of the analysis of the GW170817 binary merger event.
@@
Figure 7: Mass-Tidal Deformability (MTD) curves for neutron stars. Each curve represents a different hybrid EOS, with color coding and line styles as in fig. 2: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4; solid (dashed) lines correspond to potential 7a (5b). The top-left panel corresponds to matching procedure (b), while the others show procedure (a) with $n_{t}=\{1.2,1.5,1.8\}n_{S}$.
@@
Figure 8: Density as a function of radial distance from the center of neutron stars, for typical stars of mass $1.4M_{\odot}$. Top left panel: Matching procedure (b). Remaining panels: Matching procedure (a) with matching density $n_{t}=\{1.2,1.5,1.8\}n_{S}$. The color coding and line styles are as in fig. 2: Green for HLPS soft, orange for HLPS intermediate, red for HLPS stiff, and purple for SLy4; solid (dashed) lines correspond to potential 7a (5b). Thick lines, thin lines, and darker thick lines correspond respectively to the holographic core, the low density core (HPLS or SLy4) and the crust of the stars. Note that in the bottom-right panel, only one red curve appears; this is because for this star the density never reaches the matching density $n_{t}=0.288\rm{fm}^{-3}$, and so it is insensitive to the holographic model.
@@ 7 Conclusions
With this work, we have studied the effects of isospin breaking in nuclear matter within the holographic V-QCD model. A crucial simplifying approximation utilized in this work is the homogeneous Ansatz for nuclear matter, which consists of introducing a finite baryonic density via a discontinuity in the bulk gauge fields, which are taken to be independent of the position in three dimensional space. Isospin is introduced via an external chemical potential, following the usual holographic prescription of encoding its value in the UV boundary condition for the field $A^{a=3}_{t}$. The approximation inherent in the homogeneous Ansatz is expected to be reliable only at high densities: For this reason, we used a set of phenomenological equations of state for the description of the low-density regime (SLy4 and three representative HLPS). We considered two different approaches to matching the low density EOS with the holographic high density ones: The one labeled “procedure (a)” is more phenomenology-driven and corresponds to the introduction of a new parameter that acts as a rescaling factor for the holographic action, while with the one labeled “procedure (b)” avoid introducing new parameters and instead fit the already present ones to correctly reproduce the value of saturation density.
@@
Finally, having the computed EOS in hand, we solved the TOV equations and derived the properties of static neutron stars. We computed MR curves and tidal deformabilities of the stars resulting from each EOS, finding results that are in all cases close to observations, and a subset of EOSs that satisfy the constraints from NICER on the MR curves and from LIGO/Virgo on the tidal deformability. In tab. 1 we report the values of the parameters $n_{t},b_{1},c_{b}$ for each hybrid equation of state. On the left are the values for potentials 7a, while on the right for potentials 5b. In the last column labeled “Data” we evaluate the compatibility of the resulting neutron stars with the NICER and LIGO/Virgo data, indicating with a check mark ✓ full compatibility, while with ✗ the presence of tension with the data currently available. An interesting result is that those EOSs that produce neutron stars that satisfy these bounds are also those that correspond to parameters that correctly reproduce the phenomenological symmetry energy at saturation density: In particular, all of the hybrid EOSs built from the stiff HLPS low-density EOS are in tension with astrophysical bounds, and they all result in a high value for the symmetry energy at saturation $S_{0}=[36.4,39.4]$ MeV. For the ones built from the soft HLPS EOS we find that only one pair (corresponding to potential 5b and 7a) of hybrid EOSs is in good agreement with observational bounds, resulting from the matching at $n_{t}=1.5n_{S}$, for which $S_{0}=\{32.0,32.8\}$MeV. For the intermediate HLPS, the hybrid EOS pair that is most compatible with astrophysical bounds is that coming from the matching at $n_{t}=1.8n_{S}$, and the corresponding symmetry energy at saturation is $S_{0}=\{31.9,32.6\}$MeV. Finally, for the SLy4 EOS, both the hybrid pairs corresponding to $n_{t}={1.5,1.8}n_{S}$ are close to observational bounds: For $n_{t}=1.5n_{S}$ we have $S_{0}=\{33.7,34.3\}$MeV, while for $n_{t}=1.8n_{S}$ we find $S_{0}=\{31.3,32.0\}$MeV. This precise agreement with phenomenology may hint that the rescaling procedure (a) tends to successfully capture some physics that is lost when taking the approximations of homogeneity and of the flat-space dilaton potential.
@@
Table 1: Summary of the EOSs and parameters for matching procedure (a).
These and other improvements are for now left for future works.
Acknowledgments
Appendix A Vanishing contribution of the variation of $r_{c}$ to the symmetry energy
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{2}\mu_{I}^{2},$ (59)
$\mathcal{L}_{0}(r_{c})+\mathcal{L}_{2}(r_{c})\mu_{I}^{2}=0.$ (60)
We can then expand also the critical radius $r_{c}$ as:
$r_{c}=r_{c0}+\delta r_{c},$ (61)
with $r_{c0}$ being the critical position of the discontinuity for symmetric matter, and $\delta r_{c}$ being its $\mu_{I}$-dependent infinitesimal variation. Imposing the equilibrium condition at the unperturbed order allows us to determine $\delta r_{c}$, which is also quadratic in $\mu_{I}$: